Add without using number line:
 
(a) 11 + (–7)
(b) (–13) + (+18)
(c) (–10) + (+19)
  
yhx,kyhk,tktlyu8l;8ypy

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


(a) 11 + (-7) = 4

(b) (-13) + (+18) = 5

(c) (-10) + (+19) = 9


Thanks